A baseball is hit with a speed of 36.0 m/s . Calculate its height if it is hit at an angle of 30.0 ∘ Express your answer with the appropriate units.

A baseball is hit with a speed of 36.0 m/s . Calculate its distance traveled if it is hit at an angle of 45.0 ∘ Express your answer with the appropriate units.

A baseball is hit with a speed of 36.0 m/s . Calculate its height if it is hit at an angle of 60.0 ∘ Express your answer with the appropriate units.

A baseball is hit with a speed of 36.0 m/s . Calculate its distance traveled if it is hit at an angle of 60.0 ∘ Express your answer with the appropriate units.
Answer #1

For a projectile we have:

Maximum height reached =

Range(distance traveled) =

1) Given u = 36 m/s and = 30 o, height reached is given as:

2)Given u = 36 m/s and = 45 o, range is given as:

3)Given u = 36 m/s and = 60 o​​​​​​​, height reached:

4)Given u = 36 m/s and = 60 o​​​​​​​, so range :